# Enhancing Multimodal Large Language Models with Vision Detection Models: An Empirical Study

### v2

@@ Appendix D Model Architecture Rationale
In the main body of our paper, we exclusively select LLaVA-1.5  [^1] We do not use LLaVA-1.6 (LLaVA-NeXT) in our experiments as it is concurrent to our work, and its training code is un-accessible until the NeurIPS submission deadline. for experimentation, considering it representative of the majority of state-of-the-art models. In this section, we will illustrate this choice from the following two aspects:
@@
We choose LLaVA-1.5 as we are in a highly dynamic field and it is representative enough of most state-of-the-art MLLMs. The advanced MLLMs typically consist of three main modules: an image encoder, an input projector, and a LLM Backbone. LLaVA-1.5 adheres to this structure.
Images are first processed through the image encoder and the input projector. Most advanced MLLMs typically include a dedicated branch like this for processing image features into analogous image token sequences. Next, the image tokens are concatenated with text tokens that represent input sentences and fed into the LLM. Specifically, the text tokens representing our detection information can be directly concatenated with the extracted image tokens and used during MLLM’s training and inference. In other words, as long as the MLLM conforms to this structure, the additional detection information can be processed similarly before being inputted into the LLM, specifically being concatenated with image features at the embedding level. The way they are infused into the backbone LLM is similar. Therefore, the results of experiments conducted on LLaVA-1.5 can be applied to other MLLMs with similar structures.
Furthermore, LLaVA-1.5 has proven to be highly successful, spawning numerous outstanding works. We conduct our study based on LLaVA-1.5, enabling the application of our experimental findings to the subsequent works of LLaVA-1.5. Consequently, we are positioned to contribute to the open-source community.
@@
Specifically, we conduct visual instruction tuning on Qwen-VL (which has not undergone visual instruction tuning) with the Instruction-following Dataset of LLaVA-1.5. We compare three models: Qwen-VL-IT, Qwen-VL-IT-TFI, and Qwen-VL-IT-RBI.
  • Qwen-VL-IT refers to Qwen-VL undergoing regular visual instruction tuning. During the training and inference process, Qwen-VL-IT doesn’t incorporate detection modules.
  • Qwen-VL-IT-TFI follows the same training process as Qwen-VL-IT, but it infuses detection information during inference, corresponding to the TFI fusion strategy in our paper.
  • Qwen-VL-IT-RBI refers to fine-tuning Qwen-VL while simultaneously infusing detection information during training and inference, corresponding to the RBI fusion strategy in our paper (with a slight difference as it does not infuse detection information during the pretraining process).
@@
Table 13: Comparison between “Qwen-VL-IT”, “Qwen-VL-IT-TFI”, and “Qwen-VL-IT-RBI”.
@@
Qwen-VL-IT 79.4 77.8 49.9 1393.1 336.4 85.9 63.3 57.7 32.3 58.9
Qwen-VL-IT-TFI 78.1 80.9 49.7 1355.8 278.9 89.9 59.4 52.8 32.4 56.7
Qwen-VL-IT-RBI 78.9 82.2 54.4 1501.2 407.1 89.7 63.9 57.7 34.4 58.9
Based on Table 13, it is evident that the visual grounding capability of Qwen-VL-IT-TFI has improved compared to Qwen-VL-IT, resulting in significant score increases on the POPE benchmarks. However, Qwen-VL-IT-TFI exhibits varying degrees of decline in other tasks, similar to the results of TFI fusion strategy on the main page.
On the other hand, Qwen-VL-IT-RBI exhibits improvements across all benchmarks compared to both Qwen-VL-IT and Qwen-VL-IT-TFI, with notable score increases in both object detection and text recognition benchmarks. This mirrors the results of RBI fusion strategy on the main page, indicating that by infusing detection information during training, the model can better comprehend the detection information and consequently use it more effectively to address issues.
In summary, we elucidate the reasons behind LLaVA-1.5’s capability to serve as a representative model for many state-of-the-art MLLMs. We assert that the insights drawn from experiments on LLaVA-1.5 are broadly applicable to other structurally similar MLLMs. Furthermore, we conduct additional experiments on another MLLM, Qwen-VL, thereby demonstrating the extensive validity of our research findings.
@@ D.2 How DINO and PaddleOCRv2 Represent Other Detecion Models?
Due to the specific textual format we designed for detection information, we can process the outputs of any object detection model and OCR model into detection information, as long as they can output the names of targets, the content of texts, and the corresponding coordinates of targets. (“Here are the central coordinates of certain objects in this image: 2 people:[0.25, 0.12], [0.11, 0.43], 1 cake:[0.42, 0.32].” or “Here are the central coordinates of certain texts in this image: ‘Birthday’[0.41, 0.85], ‘YEARS’[0.11, 0.34].” ) In other words, the selection of object detection models and OCR models is not crucial. We only need to choose an excellent model based on performance and efficiency.
To better elucidate this point, we replace DINO with YOLOv8 and repeat the FTBI experiments, yielding the outcomes in Table 14.
@@
According to the table, both models bring similar performance improvements to the MLLM, suggesting that when the functionalities and performances of detection models are similar, their impact on the MLLM’s enhancement is also similar.
D.3 How Does Text-Based Detection Information Feeding Compare to Other Methods?
Besides the text-based approach, we conduct experiments on feeding the output vectors of DINO and PaddleOCRv2 into LLM, using either MLP layers or multiple cross-attention layers to map these vectors into LLM’s semantic space. However, these methods do not perform as well as the text-based approach. Below is the performance comparison among them, all of which are trained on 7B models using the RBI (Retraining Based Infusion) strategy.
Table 15: Comparison between the MLP approach, the cross-attention approach, and the text-based approach.
@@ D.3 How Does Text-Based Detection Information Feeding Compare to Other Methods?
VQAv2 VQAT MMEC POPE
MLP 77.7 57.5 268.6 86.9
Cross-attention 77.2 57.4 265.4 85.9
Text-based 78.5 60.0 412.9 89.3
It can be observed that the performance improvements introduced by DINO and PaddleOCRv2 are not manifested in the MLP method and cross-attention method. We speculate that the occurrence of this phenomenon stems from insufficient training data, which is inadequate to train the newly initialized MLP module and cross-attention module. Unlike the outputs of CLIP-ViT, the output vectors of DINO and PaddleOCRv2 have not undergone contrastive learning with text, leading to significant discrepancies between these vectors and the text feature. Hence, more data is required to train the input projector effectively. However, introducing more data to train these newly introduced input projectors would render unfair performance comparisons with LLaVA-1.5.
Nevertheless, the text-based approach can effectively transmit the output information from DINO and PaddleOCRv2 to LLM without requiring additional training data. Therefore, we use the text-based approach.

### v3

@@ Appendix D Model Architecture Rationale
On the main page of our paper, we exclusively select LLaVA-1.5 for experimentation, considering it representative of most advanced models. In this subsection, we will explain this choice from the following two aspects:
@@
We choose LLaVA-1.5 as we are in a highly dynamic field and LLaVA-1.5 is representative enough of most SOTA MLLMs. The advanced MLLMs typically consist of three main modules: an image encoder, an input projector, and a LLM backbone. LLaVA-1.5 adheres to this structure.
The process begins by encoding images into image features with an image encoder and aligning them with text features using an input projector. Most advanced MLLMs include a dedicated branch like this for processing images into analogous image token sequences. These image tokens are then combined with text tokens representing input sentences and inputted into the LLM.
Following this structure, the tokens derived from textual detection information can be directly combined with image tokens and used during MLLM’s training and inference. In other words, as long as the MLLM conforms to this structure, the additional textual detection information can be processed similarly before being inputted into the LLM and serves a similar function during training and inference. Therefore, the results of experiments conducted on LLaVA-1.5 can be applied to other MLLMs with similar structures.
Furthermore, LLaVA-1.5 has proven to be highly successful, spawning numerous outstanding works. We conduct our study based on LLaVA-1.5, enabling the application of our experimental findings to the subsequent works of LLaVA-1.5.
@@
Specifically, since the instruction-following dataset of Qwen-VL-Chat is not open-sourced, we conduct visual instruction tuning on Qwen-VL (which has not undergone visual instruction tuning) with the instruction-following dataset of LLaVA-1.5. We compare three models: Qwen-VL-IT, Qwen-VL-IT-TFI, and Qwen-VL-IT-FTBI:
  • Qwen-VL-IT refers to Qwen-VL undergoing regular visual instruction tuning. During the training and inference process, Qwen-VL-IT doesn’t infuse textual detection information.
  • Qwen-VL-IT-TFI follows the same training process as Qwen-VL-IT, but it infuses textual detection information during inference, corresponding to the TFI training strategy on the main page.
  • Qwen-VL-IT-FTBI refers to fine-tuning Qwen-VL-IT while simultaneously infusing detection information during training and inference, corresponding to the FTBI training strategy on the main page.
@@
Table 10: Comparison between “Qwen-VL-IT”, “Qwen-VL-IT-TFI”, and “Qwen-VL-IT-FTBI” on 10 well-recognized MLLM benchmark.
@@
Qwen-VL-IT 80.8 82.1 61.7 1474.8 388.9 86.5 71.5 67.5 44.7 62.9
Qwen-VL-IT-TFI 80.1 $\downarrow$ 82.5 $\uparrow$ 61.4 $\downarrow$ 1455.47 $\downarrow$ 438.9 $\uparrow$ 89.5 $\uparrow$ 69.4 $\downarrow$ 66.6 $\downarrow$ 40.3 $\downarrow$ 63.1 $\uparrow$
Qwen-VL-IT-FTBI 81.0 $\uparrow$ 82.7 $\uparrow$ 61.9 $\uparrow$ 1514.3 $\uparrow$ 417.1 $\uparrow$ 89.5 $\uparrow$ 72.9 $\uparrow$ 68.6 $\uparrow$ 46.7 $\uparrow$ 63.1 $\uparrow$
Based on Table 10, it is evident that the visual grounding capability of Qwen-VL-IT-TFI has improved compared to Qwen-VL-IT, resulting in significant score increases on the POPE benchmark and the MME-Cognition benchmark. However, Qwen-VL-IT-TFI exhibits varying degrees of decline on other tasks, similar to the results of the TFI strategy on the main page.
On the other hand, Qwen-VL-IT-FTBI exhibits comprehensive improvements across all 10 benchmarks compared to Qwen-VL-IT and Qwen-VL-IT-TFI, with notable score increases in both object detection benchmarks and text recognition benchmarks. This mirrors the results of the FTBI training strategy on the main page, indicating that by infusing textual detection information during training, the model can better comprehend the detection information and consequently use it more effectively to address issues.
Table 11: If we do not infuse detection information to Qwen-VL-IT-FTBI during inference, its performance will be on par with Qwen-VL-IT. “w/o DI” is an abbreviation for “without detection information.”
VQAv2 GQA* VQAT MMEP MMEC POPE MMB MMBCN MM-Vet SEED
Qwen-VL-IT 80.8 82.1 61.7 1474.8 388.9 86.5 71.5 67.5 44.7 62.9
Qwen-VL-IT-FTBI w/o DI 80.6 81.8 60.9 1470.9 376.4 86.6 72.0 68.3 43.9 62.5
Additionally, as shown in Table 11, we evaluate the performance of Qwen-VL-IT-FTBI without infusing detection information during inference and find that its results are comparable to those of Qwen-VL-IT. This further supports the experimental conclusion presented in the main page: fine-tuning the original MLLM allows it to retain its ability to comprehend image features derived from the image encoder, leading to strong performance on both image-level tasks and fine-grained image recognition tasks.
(3) The empirical support on LLaVA-NeXT.
we conduct the FTBI experiment again using LLaVA-NeXT, aiming to investigate whether a more advanced MLLM can enhance the performance of the FTBI model. The selected base model is llama3-llava-next-8b, and the training dataset is LLaVA-NeXT’s visual instruction tuning dataset. The results are presented as follows.
Table 12: Comparison between “LLaVA-NeXT-8B”, “LLaVA-NeXT-8B-TFI”, and “LLaVA-NeXT-8B-FTBI” on 10 well-recognized MLLM benchmark.
Model VQAv2 GQA* VQAT MMEP MMEC POPE MMB MMBCN MM-Vet SEED
LLaVA-NeXT-8B 82.7 82.8 65.1 1588.2 379.3 86.9 72.9 69.6 42.2 66.2
LLaVA-NeXT-8B-TFI 82.0 82.7 65.3 1525.9 468.9 90.3 72.0 70.8 43.8 65.5
LLaVA-NeXT-8B-FTBI 82.5 83.0 65.7 1563.9 445.0 89.4 74.0 70.3 44.1 67.0
From Table 12, incorporating detection information improves LLaVA-NeXT’s performance on benchmarks related to object detection and text recognition. Moreover, the LLaVA-NeXT version of the FTBI model demonstrates superior overall performance compared to both the original LLaVA-NeXT and the TFI model. These results align with the experimental conclusions presented on the main page.
In summary, we elucidate the reasons behind LLaVA-1.5’s capability to serve as a representative model for many advance MLLMs. We assert that the insights drawn from experiments on LLaVA-1.5 are broadly applicable to other MLLMs with similar structure. Furthermore, we conduct additional experiments on other MLLMs, Qwen-VL and LLaVA-NeXT, thereby demonstrating the extensive validity of our research findings.
@@ D.2 How DINO and PaddleOCRv2 Represent Other Detecion Models?
Due to the specific textual format we designed, we can process the outputs of any object detection models and OCR models into textual detection information, as long as they can output the names of targets, the content of texts, and the corresponding coordinates of targets. (“Here are the central coordinates of certain objects in this image: 2 people:[0.25, 0.12], [0.11, 0.43], 1 cake:[0.42, 0.32].” or “Here are the central coordinates of certain texts in this image: ‘Birthday’[0.41, 0.85], ‘YEARS’[0.11, 0.34].” ) In other words, the selection of object detection models and OCR models is not crucial. We can choose any detection models for the experiments.
@@
To better elucidate this point, we replace DINO with another object detection model, YOLOv8, and repeat the FTBI experiments, yielding the outcomes in Table 13. According to the table, both models bring similar performance improvements to the studied MLLM, suggesting that when the functionalities and performances of detection models are similar, their impact on the MLLM’s enhancement is also similar.
D.3 Experiments on Detectors with Varying Performance
The outputs of low-performance detection models often include noise, which can adversely affect the following MLLM. To investigate the impact of detection model accuracy on the MLLM performance, we employ a low-performance detection model YOLOv5N (Jocher et al., 2023) (mAP 34.3) and a high-performance detection model YOLOv11L (mAP 53.4) (replacing only the object detection model DINO while keeping the PaddleOCR unchanged), conduct both the training-free and fine-tuning experiments again and compare the performance gains brought by them. The results are presented in Table 14.
Table 14: Experiments based on YOLOv5N and YOLOv11L.
@@ D.3 Experiments on Detectors with Varying Performance
Model VQAv2 GQA* VQAT MMEP MMEC POPE MMB MMBCN MM-Vet SEED
LLaVA-1.5-7B 78.5 79.6 58.2 1510.7 355.7 85.9 64.3 58.3 30.5 58.6
LLaVA-1.5-7B-YOLOv5N-TFI 78.3 79.3 59.0 1459.9 382.9 86.3 64.2 56.3 32.2 59.9
LLaVA-1.5-7B-YOLOv5N-FTBI 78.6 79.9 60.0 1492.7 402.1 87.1 68.9 62.5 33.5 60.4
LLaVA-1.5-7B-YOLOv11L-TFI 78.5 79.5 59.0 1490.6 364.6 87.9 64.7 56.5 33.8 60.3
LLaVA-1.5-7B-YOLOv11L-FTBI 79.0 80.0 60.2 1497.5 405.4 88.9 70.3 62.9 34.6 60.6
The results are presented in the table, from which the following conclusions can be drawn:
  • Under the training-free strategy, YOLOv5N introduces noise to LLaVA-1.5-7B, resulting in performance degradation. In contrast, YOLOv11L, due to its superior performance, introduces minimal noise, thereby causing negligible negative impact.
  • For object detection-related tasks (POPE & MM-Vet), both YOLOv5N and YOLOv11L contribute to performance improvements under the training-free strategy. However, the improvement achieved by YOLOv5N is evidently smaller than that of YOLOv11L, which can be attributed to the disparity in their detection capabilities. This highlights the training-free strategy’s limited adaptability to low-performance detection models.
  • Furthermore, after fine-tuning, both two versions of the MLLM achieve comprehensive performance improvements, surpassing the original LLaVA-1.5-7B. The results align with the conclusions on the main page, demonstrating that the fine-tuning strategy enables the MLLM to better differentiate between noise and useful information and more effectively interpret specially designed detection information, leading to performance enhancement.
These results indicate that the fine-tuning strategy is more robust and better able to handle the erroneous information introduced by low-performance detection models compared to the training-free strategy.
D.4 Model Fine-Tuning Without the Use of Detection Data
On the main page, the fine-tuning dataset we used includes object detection data. In this subsection, we will explore fine-tuning the MLLM using data unrelated to detection tasks and examine whether the FTBI model can still retain its good language understanding capabilities.
Regarding the new fine-tuning dataset, we remove samples related to “coordinate” questions (object detection samples) and eliminate all text recognition samples from the original LLaVA fine-tuning dataset. Consequently, the number of samples decreases from 665K to 450K. The experimental results are presented in the table below, and the corresponding model name is ”LLaVA-1.5-7B-FTBI-FNDI”.
Table 15: Results of fine-Tuning the model without using detection data.
Model VQAv2 GQA* VQAT MMEP MMEC POPE MMB MMBCN MM-Vet SEED
LLaVA-1.5-7B 78.5 79.6 58.2 1510.7 355.7 85.9 64.3 58.3 30.5 58.6
LLaVA-1.5-7B-TFI 78.5 79.2 59.2 1497.0 401.0 89.9 65.0 57.2 33.7 60.6
LLaVA-1.5-7B-FTBI-FNDI 79.1 79.8 59.5 1518.0 410.4 88.8 68.4 60.3 33.9 61.1
LLaVA-1.5-7B-FTBI 79.0 80.1 60.1 1482.7 397.9 88.9 67.3 60.2 35.2 60.8
From Table 15, it is evident that even without fine-tuning on detection-related data, the FTBI model still demonstrates strong performance, significantly surpassing the original model and the training-free model. Moreover, its results are only slightly below the version fine-tuned with detection data. These results indicate that, even without fine-tuning on tasks related to detection, the fine-tuned model is still capable of maintaining a broad range of language understanding abilities.
D.5 Model Fine-Tuning With an Unfrozen Visual Encoder
On the main page, we do not train the visual encoder because the baseline we use, LLaVA-1.5-7B, also keeps the visual encoder frozen during training. In this subsection, we unfreeze the visual encoder and repeat both the retraining and fine-tuning processes for exploration. The results are presented as follows, where “TVE” denotes training with the visual encoder unfrozen.
Table 16: Results of training with the visual encoder unfrozen.
Model VQAv2 GQA* VQAT MMEP MMEC POPE MMB MMBCN MM-Vet SEED
LLaVA-1.5-7B 78.5 79.6 58.2 1510.7 355.7 85.9 64.3 58.3 30.5 58.6
LLaVA-1.5-7B-TFI 78.5 79.2 59.2 1497.0 401.0 89.9 65.0 57.2 33.7 60.6
LLaVA-1.5-7B-RBI-TVE 78.2 76.1 59.3 1466.5 396.4 89.1 67.2 60.4 34.0 60.5
LLaVA-1.5-7B-FTBI-TVE 79 79.7 60.4 1556.9 412.1 89.3 68.9 61.2 34.6 60.8
Table 17: Results of training with the visual encoder unfrozen (without detection information being input during inference).
[…]
LLaVA-1.5-7B 78.5 79.6 58.2 1510.7 355.7 85.9 64.3 58.3 30.5 58.6
LLaVA-1.5-7B-RBI-TVE w/o DI 76.4 75.4 56.1 1480.7 289.3 83.1 66.3 59.5 30.1 59.6
LLaVA-1.5-7B-FTBI-TVE w/o DI 78.1 78.9 57.7 1499.6 318.6 85.5 66.8 60.1 30.8 60.5
As shown in Table 16, even with the visual encoder being trained, the performance of the training-free, retraining, and fine-tuning strategies aligns with the patterns summarized on the main page. Specifically, the RBI model outperforms the training-free model, while the FTBI model further surpasses the RBI model. Moreover, the fine-tuned model achieves the best performance in 9 out of 10 benchmarks while training with the visual encoder unfrozen.
Furthermore, Table 17 presents the performance of RBI and FTBI models when the detection information is not dynamically input during inference. It demonstrates that, under the condition where the visual encoder is unfrozen, the fine-tuned model still maintains comparable performance to the original LLaVA-1.5-7B, while the RBI model performs worse than the original model. This indicates that the fine-tuning strategy better balances the contributions of the visual encoder’s outputs and the detection information, thereby facilitating a more effective understanding of detection cues. These findings are consistent with the conclusions presented in our paper.
D.6 Experiments on a Detector with Broader Detection Ranges
On the main page, the object detection model we use, DINO, is limited to detecting 80 object categories, as it is trained on the MS-COCO (Lin et al., 2014) dataset. In this subsection, we explore whether using an object detection model with a broader detection range could further improve the performance of the FTBI model. To this end, we select Co-DETR-LVIS (Zong et al., 2023), which is trained on the LVIS (Gupta et al., 2019) dataset and can detect 1,203 object categories. We conduct both training-free and fine-tuning experiments using Co-DETR-LVIS, and the results are as follows:
Table 18: Experimental results based on Co-DETR-LVIS.
Model VQAv2 GQA* VQAT MMEP MMEC POPE MMB MMBCN MM-Vet SEED
LLaVA-1.5-7B 78.5 79.6 58.2 1510.7 355.7 85.9 64.3 58.3 30.5 58.6
LLaVA-1.5-7B-DINO-TFI 78.5 79.2 59.2 1497.0 401.0 89.9 65.0 57.2 33.7 60.6
LLaVA-1.5-7B-DINO-FTBI 79.0 80.1 60.1 1482.7 397.9 88.9 67.3 60.2 35.2 60.8
LLaVA-1.5-7B-CoDETR-LVIS-TFI 77.7 76.9 58.5 1465.4 386.8 87.4 65.7 57.3 33.9 60.1
LLaVA-1.5-7B-CoDETR-LVIS-FTBI 78.7 79.5 59.7 1469.1 387.1 88.4 66.6 60.1 35.6 60.7
We can derive the following points from the table:
  • Under the training-free condition, the TFI model based on Co-DETR-LVIS performs worse than the DINO-based TFI model across almost all benchmarks. After analysis, we believe that this is because Co-DETR-LVIS introduces more noise compared to DINO, as it detects a significant number of redundant objects.
  • After fine-tuning, the MLLM gains the ability to mitigate the noise introduced by Co-DETR-LVIS. Consequently, the FTBI model based on Co-DETR-LVIS achieves comprehensive performance improvements over its TFI counterpart. This observation is consistent with the conclusions presented in our paper.
  • Furthermore, when comparing the FTBI model based on Co-DETR-LVIS with the FTBI model based on DINO, it is evident that the Co-DETR-LVIS-based model performs worse, exhibiting inferior results across all ten benchmarks.
In summary, detection models with a wider range of object categories do not necessarily improve the performance of the FTBI models. We think this is because many of the objects they detect are redundant and may instead introduce noise, leading to a decrease in performance scores.
D.7 Further Discussion on Related Works
(1) Why we conduct comparative experiments around adaptive training based on “deploying detection models to assist MLLMs”?
Deploying independent detection models (or models for other downstream tasks) to generate auxiliary text for MLLMs is both straightforward and effective. By simply incorporating external text descriptions into the MLLMs, it significantly improves their performance. Moreover, the deployed models are interchangeable, allowing for convenient updates and the replacement with higher-performing models, thereby enhancing the overall performance of the framework. Given its numerous advantages, an increasing number of researchers are investigating this paradigm and working based on it.
Nevertheless, many researchers tend to adopt training-free strategies. The impact of adaptive training, however, remains an important area of investigation. Therefore, we conduct systematic experiments based on the training-free and adaptive training strategies in this paradigm, as there has not been a comprehensive comparison between them.
(2) Distinctions from approaches involving the introduction of special tokens.
In the academic community, there is a paradigm also focusing on detection information, which involves introducing special tokens to explicitly infuse detection information in both input and output, guiding MLLMs to leverage this information. Typical methods include MiniGPT4-v2 (Chen et al., 2023b), VisionLLM (Wang et al., 2024b), and Shikra (Chen et al., 2023c).
Nevertheless, this paradigm differs significantly from the paradigm we focus on.
  • First, the method of deploying detection models allows MLLMs to receive real-time detection information during both training and inference. This type of detection information encompasses the locations of all detectable objects in the image, containing rich details about the image. In contrast, the special token method, which does not deploy detection models, requires manual input of detection information at the input stage. Such detection information is typically limited to a single object or a small number of objects, serving primarily as task guidance. Thus, the role of detection information differs between these approaches: in the former, it assists MLLMs for downstream tasks by providing useful detection details, while in the latter, it usually serves only as a signal, indicating that the task involves detecting specific targets.
  • Furthermore, the detection information introduced by MiniGPT4-v2 and VisionLLM is completely accurate, as it is derived from datasets. In contrast, deployed detection models may occasionally produce errors, introducing noise that affects the training-free model. This noise, however, also trains the MLLMs’ ability to denoise.
Therefore, the focus of our paper is fundamentally different from them. The training strategies for deploying detection models to assist MLLMs have not been as extensively explored as methods involving the special tokens. Our systematic study on this topic represents a new departure.
(3) Our study is a pioneering work, offering inspiration for further research.
Our research investigates whether adaptive training can help MLLMs better identify noise in real-time detection information and more effectively leverage the outputs of additional detection models to enhance VQA performance. To the best of our knowledge, no previous work has systematically explored the impact of adaptive training on deploying detection models to assist MLLMs. To draw inspiration from it, we conduct a series of systematic experiments in this direction.
Our findings demonstrate that the adaptive training strategy indeed outperforms the training-free strategy. Additionally, we confirm that fine-tuning with only a small amount of high-quality VQA data can also lead to improved performance, and the performance gain is still preserved even after replacing the detection models. Similar empirical evidences also presented in several subsequent works Chen et al. (2024c); Jiao et al. (2024).
As a pioneering study in this area, we have uncovered many valuable insights, and we hope our findings can provide insights for researchers in the relevant field.